# DexSinGrasp: Learning a Unified Policy for Dexterous Object Singulation and Grasping in Cluttered Environments

### v1

@@ Abstract
Grasping objects in cluttered environments remains a fundamental yet challenging problem in robotic manipulation. While prior works have explored learning-based synergies between pushing and grasping for two-fingered grippers, few have leveraged the high degrees of freedom (DoF) in dexterous hands to perform efficient singulation for grasping in cluttered settings. In this work, we introduce DexSinGrasp, a unified policy for dexterous object singulation and grasping. DexSinGrasp enables high-dexterity object singulation to facilitate grasping, significantly improving efficiency and effectiveness in cluttered environments. We incorporate clutter arrangement curriculum learning to enhance success rates and generalization across diverse clutter conditions, while policy distillation enables a deployable vision-based grasping strategy. To evaluate our approach, we introduce a set of cluttered grasping tasks with varying object arrangements and occlusion levels. Experimental results show that our method outperforms baselines in both efficiency and grasping success rate, particularly in dense clutter. Codes, appendix, and videos are available on our project website https://nus-lins-lab.github.io/dexsingweb/.
@@ I Introduction
One approach to handling densely cluttered environments is to singulate the target object from surrounding objects. Researchers have explored frameworks to learn the synergies between pushing and grasping [12, 13, 14] for two-fingered grippers. These methods emphasize that in densely cluttered scenes where the target object is initially ungraspable, singulation must first be performed before grasping [15, 16, 17, 18, 19, 20, 21, 22, 23]. Naturally, this leads to a pushing-and-grasping process, where Zeng et al. [12] proposed a method using two networks to coordinate pushing and grasping. However, their approach relies solely on a grasp success reward to guide coordination, leading to low sample efficiency during training. Other methods introduced adversarial training to enhance synergy learning efficiency[13, 14]. Despite these advancements, the mechanical constraints of grippers require the target object to be fully isolated from surrounding clutter, making singulation inefficient. In contrast, dexterous hands perform singulation using only their fingers, minimizing movement of the end-effector (i.e., the palm) and providing a more flexible and efficient approach to object rearrangement in cluttered settings.
Figure 1: We propose DexSinGrasp to learn a unified policy for dexterous object singulation and grasping in cluttered environments
@@
Learning singulation for grasping with dexterous hands offers a promising approach to handling objects in cluttered environments. However, the high degrees of freedom (DoF) of dexterous hands and the complexity of cluttered scenes make this synergy challenging to learn. One approach to addressing this challenge is task decomposition. Jiang et al. [24] propose a three-stage framework where separate policies are trained for isolating, grasping, and lifting. While this decomposition simplifies learning by breaking the problem into manageable sub-tasks, it also limits the synergy between singulation and grasping. In particular, minimal singulation may often suffice for dexterous grasping, as illustrated in Fig. 1. Alternatively, curriculum learning [25] has proven effective in tackling complex tasks and has already been successfully applied to dexterous grasping policies [26, 8]. Our work builds on this success by exploring curriculum learning to develop a unified policy that integrates both singulation and grasping for dexterous manipulation.
In this work, we develop a reinforcement learning framework to train a unified policy that seamlessly integrates object singulation and grasping. This framework enables a dexterous hand to efficiently grasp target objects from tightly cluttered environments, as illustrated in Fig. 1. Due to the challenges of directly solving grasping tasks in general cluttered environments, our method leverages clutter arrangement curriculum learning to progressively enhance the performance of the teacher policy in generated cluttered environments with increasing complexity in object quantity, types, and arrangements. Furthermore, through teacher-student policy distillation, we obtain a vision-based student policy that generalizes across diverse cluttered environments and can be deployed on a real-world robot.
  • We develop a unified reinforcement learning policy for dexterous object singulation and grasping, enabling dexterous hands to effectively and efficiently grasp objects in tightly cluttered environments.
  • We incorporate clutter arrangement curriculum learning to improve policy performance across various cluttered scenes and employ policy distillation to obtain a vision-based grasping policy suitable for real-world deployment.
  • We design a set of cluttered grasping tasks and experiments with varying difficulty levels and conduct extensive experiments to demonstrate the effectiveness and efficiency of our proposed DexSinGrasp.
Our method leverages a unified reward design to transform target objects—initially ungraspable due to tight occlusions—into graspable states, ultimately completing the grasp.
@@ II Related Work
Dexterous Grasping Policy. Recent advancements in dexterous grasping policies leverage reinforcement learning (RL) and physics simulations to model dynamic hand-object interactions. Due to the learning challenges introduced by high DoF, some approaches decompose grasping into hierarchical levels or unify multiple grasping objectives [4, 5, 6, 7] to improve scalability. In contrast, others focus on obtaining deployable vision-based policies for real-world applications through policy distillation [8, 9]. Recently, dexterous grasping in cluttered environments has gained increasing attention. Through encouraging visual exposure of the target object [10], and extracting scene information from segmented point clouds [10] or segmented images [11], grasping in relatively loose clutter conditions is successfully achieved. However, rather than solely focusing on grasping, our approach learns to perform object singulation for grasping, leading to a significantly higher success rate due to the sufficient grasping space, especially in tightly cluttered environments.
Grasping in Object Clutters. Prior studies on robotic grasping in cluttered environments with grippers have explored various approaches to address the challenges such as occlusions [15, 16, 17] and limited grasping space [18, 19, 20, 21, 22]. A common requirement in these scenarios is decluttering or singulation [12, 23, 14, 27, 28, 29, 30, 31, 32, 33, 13, 34], which isolates the target object to facilitate successful grasping. Zeng et al. [12] utilized two separate networks to learn the synergy between pushing and grasping with a parallel gripper, inspiring subsequent studies on the adversarial training of these two networks [13, 14]. To overcome the dexterity constraints of grippers, SOPE [24] uses fingers to isolate the target object as one of the pre-defined stages. However, our method uses clutter arrangement curriculum learning to obtain a synergistic dexterous singulation that efficiently contributes to grasping.
@@ III Method
Problem Formulation We formulate dexterous grasping and object singulation as a reinforcement learning task. Specifically, we consider a tabletop scene with the target object $b^{\text{target}}$ and $n$ surrounded objects $\left\{b_{i}\right\}_{i=1}^{n}$. This setup represents common cluttered scenarios and introduces significant grasping challenges due to the tight arrangement of surrounding objects. The target object is positioned at the center of the clutter, making direct grasping difficult. To overcome this challenge, we train robots to singulate the target object from its surroundings, thereby creating sufficient space for inserting fingers during dexterous grasping. In this work, we propose DexSinGrasp, a unified policy that jointly optimizes object singulation and dexterous grasping using reinforcement learning.
@@
where the proprioceptive robot state $s_{t}^{R}\in\mathbb{R}^{72}$ includes the wrist pose as well as joint positions, velocities, and forces for each finger and wrist dummy joints; the action $a_{t-1}$ at the previous time step will be discussed later; the object state $s_{t}^{O}\in\mathbb{R}^{16}$ consists of the object’s position and quaternion, linear and angular velocity, and object-hand position difference; the hand-object distances $d_{t}^{HO}\in\mathbb{R}^{21}$ present the minimum distances between each hand links and points on the object; the time encoding $T_{t}\in\mathbb{R}^{29}$ encodes the current time along with a sine-cosine time embedding. The singulation distance $d_{t}^{S}\in\mathbb{R}^{8}$ presents the distances between the target object and surrounding objects, indicating the level of enclosure within the clutter. If the number of surrounding objects satisfies $n<8$, the corresponding dimensions are padded with 0.
@@
Overview. As illustrated in Fig. 2, we train a unified policy for dexterous object singulation and grasping through a structured learning framework, and adopt teacher-student policy distillation for real-world deployment. First, we introduce a unified reward design (Sec. III-A) that seamlessly integrates singulation and grasping into a single objective, enabling more efficient policy learning. Then, to improve learning efficiency in cluttered environments, we adopt Clutter Arrangement Curriculum Learning (Sec. III-B) to progressively train state-based teacher policies. Finally, we employ Teacher-Student Policy Distillation (Sec. III-C) to transfer knowledge from the teacher policies to a vision-based student policy, allowing deployment on a real robot by mapping high-dimensional visual observations to effective actions.
@@
$r^{P}$ $-2.0\times d^{P}$ Palm reward to encourage hand approaching the target object.
@@
$r^{J}$ $-d^{J}$ Joint reward to encourage the finger to grasp the target object.
$r^{F}$ $-\sum_{j=1}^{h}\min_{i=1}^{p}\|p_{j}^{\text{fingertip}}-p_{i}^{\text{target}}\|_{2}$ Fingertip reward to encourage fingertip to grasp the target object.
@@ III-A Unifying Dexterous Object Singulation and Grasping
Training a policy for dexterous grasping from scratch in cluttered environments poses significant challenges due to the high-dimensional action space and complex contact interactions. The naive approach is to directly train a policy that only rewards successful grasps in cluttered environments. However, this often struggles with sample efficiency, as the policy simultaneously learns to singulate and grasp effectively. Another alternative is a two-stage method, where a singulation policy is trained separately to first singulate the object before a grasping policy is trained. However, this approach is inefficient in execution, as it fails to fully exploit the synergy between singulation and grasping, leading to longer task completion time and suboptimal transitions between the two stages.
@@
Detailed definitions of these rewards can be found in Tab. I, where $t$ subscript is omitted for simplicity. In the rewards, $r_{t}^{P}$ encourages the hand palm to stay close to the target object; $r_{t}^{J}$ and $r_{t}^{F}$ both encourage the hand to grasp the target object; $r_{t}^{L}$ encourage the hand to lift the target object once contact is made, $r_{t}^{S}$ encourages separation of the target object from the obstacles, $r_{t}^{G}$ encourages the hand to move to the goal position; $r_{t}^{B}$ is a bonus term for a successful singulation and grasping process; $d_{t}^{P}$ is the minimum distance between hand palm and target object; $d_{t}^{J}$ is the minimum distance between hand links and target object. $\left\{p_{i}^{\text{target}}\right\}_{i=1}^{p}$ are the positions of $p$ points on the target object; $p^{\text{palm}}$ is the palm position of the hand; $\left\{p_{j}^{\text{link}}\right\}_{j=1}^{m}$ are the positions of $m$ links of the hand; $\left\{p_{j}^{\text{fingertip}}\right\}_{j=1}^{h}$ are the positions of $h$ fingertips of the hand; $p^{\text{goal}}$ and $p^{\text{target}}$ are the goal position and current position for the target object; $\left\{p_{i}\right\}_{i=1}^{n}$ are the positions of $n$ obstacles; $a^{P_{t_{z}}}$ refers to the palm translation in the $+z$ direction, which corresponds to the lifting motion of the target object.
The reward function consists of two main components: an approach reward that encourages the hand to move toward the target object and a lifting reward that promotes object elevation after contact is established. To further facilitate singulation, the singulation reward $r_{t}^{S}$ is incorporated into both components, incentivizing the hand to separate the target object from surrounding obstacles. The transition between these two reward stages is achieved by a contact criterion specified by $d_{t}^{P}<0.06$ and $d_{t}^{J}<0.2$. These values are selected such that the palm and fingers are close enough for a successful grasping. The reward transitions from approaching to grasping, ensuring a smooth progression toward a successful grasp.
Despite this unified learning framework, training remains highly challenging due to the increasing complexity of cluttered environments. To further improve learning efficiency and policy generalization, we introduce Clutter Arrangement Curriculum Learning, which progressively increases clutter complexity. This structured training approach enables the teacher policy to avoid the inefficiency of multi-stage methods.
@@ III-B Clutter Arrangement Curriculum Learning
To ensure successful and efficient object singulation and grasping in tightly cluttered environments, we begin by training the teacher policy with privileged information from the simulation and adopt clutter arrangement curriculum learning, allowing our teacher policy to progressively improve as object diversity and spatial complexity increase. Using Proximal Policy Optimization (PPO) [35], we optimize the policy to maximize the cumulative discounted reward $E[\sum_{t=1}^{T}\gamma^{t-1}r_{t}]$, enabling effective reinforcement learning in cluttered environments.
III-B1 Cluttered Environment Generation
To facilitate the training and testing of policies across various object configurations, we introduce a cluttered environment generation module designed to create diverse object-based tasks. Our cluttered environment primarily consists of block-shaped objects with varying quantities (from 4 to 8) and shapes (1×1, 1×2, and 1×3 blocks). Based on the degree of enclosure, the tasks are generally divided into two categories, as shown in Fig. 3.
Dense Arrangements. This type of task arranges different quantities of the surrounding 1×1 blocks densely near the target object to create an extreme scenario that challenges the singulation and grasping policies under dense and narrow conditions.
Random Arrangements. This type of task arranges objects of different quantities and shapes randomly around the target object for grasping, mainly to test the generalization of the singulation and grasping policies.
For simplicity, we use D/R-$n$ to denote task setting with $n$ objects for dense (D) or random (R) arrangements, such as D-8.
Figure 3: Dense and random arrangement settings. We introduce a cluttered environment generation module to create diverse object settings, including different obstacle quantities from 0 to 8, dense and random arrangements, various poses, and block shapes. For simplicity, we only show obstacles with numbers 4 to 8.

### v3

# DexSinGrasp: Learning a Unified Policy for Dexterous Object Singulation and Grasping in Densely Cluttered Environments

@@ Abstract
Grasping objects in cluttered environments remains a fundamental yet challenging problem in robotic manipulation. While prior works have explored learning-based synergies between pushing and grasping for two-fingered grippers, few have leveraged the high degrees of freedom (DoF) in dexterous hands to perform efficient singulation for grasping in cluttered settings. In this work, we introduce DexSinGrasp, a unified policy for dexterous object singulation and grasping. DexSinGrasp enables high-dexterity object singulation to facilitate grasping, significantly improving efficiency and effectiveness in cluttered environments. We incorporate clutter arrangement curriculum learning to enhance success rates and generalization across diverse clutter conditions, while policy distillation enables a deployable vision-based grasping strategy. To evaluate our approach, we introduce a set of cluttered grasping tasks with varying object arrangements and occlusion levels. Experimental results show that our method outperforms baselines in both efficiency and grasping success rate, particularly in dense clutter. Codes, appendix, and videos are available on our website  https://nus-lins-lab.github.io/dexsingweb/.
@@ I Introduction
One approach to handling densely cluttered environments is to singulate—spatially separate—the target object from surrounding objects. Researchers have explored frameworks to learn the synergies between pushing and grasping [12, 13, 14] for two-fingered grippers. These methods emphasize that in densely cluttered scenes where the target object is initially ungraspable, singulation must first be performed before grasping [15, 16, 17, 18, 19, 20, 21, 22, 23]. Naturally, this leads to a pushing-and-grasping process, where Zeng et al. [12] proposed a method using two networks to coordinate pushing and grasping. However, their approach relies solely on a grasp success reward to guide coordination, leading to low sample efficiency during training. Other methods introduced adversarial training to enhance synergy learning efficiency[13, 14]. Despite these advancements, the mechanical constraints of grippers require the target object to be fully isolated from surrounding clutter, making singulation inefficient. In contrast, dexterous hands perform singulation using only their fingers, minimizing movement of the end-effector (i.e., the palm) and providing a more flexible and efficient execution in cluttered settings.
Figure 1: We propose DexSinGrasp to learn a unified policy for dexterous object singulation and grasping in densely cluttered environments
@@
Learning singulation for grasping with dexterous hands offers a promising approach to handling objects in cluttered environments. However, the high degrees of freedom (DoF) of dexterous hands and the complexity of cluttered scenes make this synergy challenging to learn. One approach to addressing this challenge is task decomposition. Jiang et al. [24] propose a three-stage framework where separate policies are trained for isolating, grasping, and lifting. While this decomposition simplifies learning by breaking the problem into manageable sub-tasks, it also limits the synergy between singulation and grasping. In particular, minimal singulation may often suffice for dexterous grasping, as illustrated in Fig. 1. Alternatively, curriculum learning [25] has proven effective in tackling complex tasks and has already been successfully applied to dexterous grasping policies [26, 8]. More recently, Bai et al. [27] utilize a dexterous hand to expose the target object from clutter for grasping, but their reward design primarily encourages palm movement during singulation. In contrast, we argue that finger dexterity should be leveraged throughout the singulation to improve efficiency. For instance, in a top-down grasp, the palm can remain above the target while the fingers actively separate surrounding objects until a feasible grasp is possible. Therefore, our work builds on this success by exploring curriculum learning to develop a unified policy that integrates both singulation and grasping with finger dexterity.
In this work, the framework enables a dexterous hand to efficiently retrieve the target object from tightly cluttered environments, as illustrated in Fig. 1. We develop a reinforcement learning (RL) framework to seamlessly integrate object singulation and grasping through a unified reward design. Due to the challenges of directly solving grasping tasks in densely cluttered environments, our method leverages clutter arrangement curriculum learning to progressively enhance the performance of the teacher policy in generated cluttered environments with increasing complexity in object quantity, types, and arrangements. Furthermore, through teacher-student policy distillation, we obtain a vision-based student policy that generalizes across diverse cluttered environments and can be deployed on a real-world robot.
  • We develop a unified RL policy for dexterous object singulation and grasping, enabling dexterous hands to effectively and efficiently retrieve objects in tightly cluttered environments.
  • We incorporate clutter arrangement curriculum learning to improve policy performance across various cluttered scenes and employ policy distillation to obtain a vision-based policy suitable for real-world deployment.
  • We design a set of cluttered grasping tasks and experiments with varying difficulty levels and conduct extensive experiments to demonstrate the effectiveness and efficiency of our proposed DexSinGrasp, and to verify the significance of finger dexterity.
@@ II Related Work
Dexterous Grasping Policy. Dexterous grasping with high degrees of freedom (DoF) presents significant learning challenges. To address these challenges, some approaches decompose grasping into hierarchical levels and use large grasping dataset to train high-level grasp proposal module [4, 5, 6, 7] in a supervised setting, followed by motion planning or low-level policies. Modern approaches on dexterous grasping benifit from RL and massive parallel simulations for dynamic hand–object interactions. Building on this foundation, others focus on obtaining deployable vision-based policies for real-world applications through policy distillation [8, 9]. Researchers have also explored population-based training for reward-scale tuning, which can significantly reduce the cost of manual hyperparameter tuning [28]. Recently, dexterous grasping in cluttered environments has gained increasing attention. Through encouraging visual exposure of the target object [10], and extracting scene information from segmented point clouds [10] or segmented images [11], grasping in relatively loose clutter conditions is successfully achieved. However, rather than solely focusing on grasping, our approach learns to perform object singulation for grasping, leading to a significantly higher success rate, especially in tightly cluttered environments.
Grasping in Object Clutters. Prior studies on robotic grasping in cluttered environments with grippers have explored various approaches to address the challenges such as occlusions [15, 16, 17] and limited grasping space [18, 19, 20, 21, 22, 29]. A common requirement in these scenarios is decluttering or singulation [12, 23, 14, 30, 31, 32, 33, 34, 35, 36, 13, 37], which isolates the target object to facilitate successful grasping. Zeng et al. [12] utilized two separate networks to learn the synergy between pushing and grasping with a parallel gripper, inspiring subsequent studies on the adversarial training of these two networks [13, 14]. Chen et al. [38] optimize contact and object trajectories and solve hand motion via inverse kinematics in reverse order, forming an open-loop control strategy, while our learning-based approach achieves adaptive closed-loop control, making it more effective in dynamic and contact-rich tasks. Due to large space needed for palm-driven pushing and limited efficiency of parallel grippers, SOPE [24] uses fingers of a dexterous hand to isolate the target object as one of the pre-defined stages. However, our method uses clutter arrangement curriculum learning to obtain a synergistic dexterous singulation that greatly improves grasping efficiency. Recent work [27] has explored using dexterous hands to expose target objects from clutter for grasping, but primarily relies on palm-driven motion during singulation. In contrast, we highlight the importance of using finger dexterity throughout the singulation to improve efficiency through our unified reward design.
@@ III Method
Problem Formulation We first define a Singulation-Necessary Clutter (SNC) as a clutter configuration in which the target object is initially ungraspable and singulation is required before any feasible grasp can be executed.
To construct SNC scenes in a controlled and reproducible way, we impose the following three sufficient conditions: (1) The target object is fully enclosed by non-target objects and located at the geometric center of the clutter; (2) The horizontal distance between the target and each surrounding object is smaller than the fingertip size; (3) The target and surrounding objects have identical height.
These constraints ensure that direct grasping is physically infeasible, and that singulation is required to create graspable space. Based on the conditions, we consider a 2D cluttered scene on a horizontal tabletop, where the target object lies in the same plane as surrounding distractor objects, with no vertical occlusion. The densely cluttered setting mentioned in the Sec. I serves as a concrete example of SNC. We further study two representative types of SNC:
Dense Arrangement. Objects are tightly packed with minimal spatial variability due to strong geometric constraints.
Random Arrangement. Objects are more loosely placed with slightly increased inter-object distances, resulting in more diverse spatial configurations.
Throughout training, we use cuboid-shaped objects as the standard shape for training. This choice facilitates the design of our curriculum learning scheme by providing consistent and controllable geometric structure. Moreover, we empirically find that policies trained on cuboids exhibit robust generalization to unseen shapes during later evaluations.
We then formulate dexterous grasping and object singulation as a RL task. Specifically, we consider a tabletop scene with the target object $b^{\text{target}}$ and $n$ surrounded objects $\left\{b_{i}\right\}_{i=1}^{n}$. To overcome this challenge, we train a policy to singulate the target object from its surroundings, thereby creating sufficient space for inserting fingers during dexterous grasping. Once singulation and grasping is complete, the target object is lifted to a goal position above the table. As a solution, we propose DexSinGrasp, a unified policy that jointly optimizes dexterous object singulation and grasping using RL.
@@
where the proprioceptive robot state $s_{t}^{R}\in\mathbb{R}^{72}$ includes the wrist pose as well as joint positions, velocities, and forces for each finger and wrist dummy joints (All joint-related dimensions hereafter are based on the Leap Hand [39] and omitted for brevity); the action $a_{t-1}$ at the previous time step will be discussed later; the object state $s_{t}^{O}\in\mathbb{R}^{16}$ consists of the object’s position and quaternion, linear and angular velocity, and object-hand position difference; the hand-object distances $d_{t}^{HO}\in\mathbb{R}^{21}$ present the minimum distances between each hand links and points on the object; the time encoding $T_{t}\in\mathbb{R}^{29}$ encodes the current time along with a sine-cosine time embedding. The singulation distance $d_{t}^{S}\in\mathbb{R}^{8}$ presents the distances between the target object and surrounding objects, indicating the level of enclosure within the clutter. If the number of surrounding objects satisfies $n<8$, the corresponding dimensions are padded with 0.
@@
Overview. As illustrated in Fig. 2, we train a unified policy for dexterous object singulation and grasping through a structured learning framework, and adopt teacher-student policy distillation for real-world deployment. First, we introduce a unified reward design (Sec. III-A) that seamlessly integrates singulation and grasping into a single objective, enabling more efficient policy learning. Then, to improve learning efficiency in SNCs, we adopt Clutter Arrangement Curriculum Learning (Sec. III-B) to progressively train state-based teacher policies. Finally, we employ Teacher-Student Policy Distillation (Sec. III-C) to transfer knowledge from the teacher policies to a vision-based student policy, allowing deployment on a real robot by mapping high-dimensional visual observations to effective actions.
@@
$r^{P}$ $2.0\times d^{P}$ Palm reward to encourage hand approaching the target object.
@@
$r^{J}$ $d^{J}$ Joint reward to encourage the finger to grasp the target object.
$r^{F}$ $\sum_{j=1}^{h}\min_{i=1}^{p}\|p_{j}^{\text{fingertip}}-p_{i}^{\text{target}}\|_{2}$ Fingertip reward to encourage fingertip to grasp the target object.
@@ III-A Unifying Dexterous Object Singulation and Grasping
Training a policy for dexterous grasping from scratch in SNCs poses significant challenges due to the high-dimensional action space and complex contact interactions. A naive approach trains a policy with rewards only for successful grasps in SNCs, but this suffers from low sample efficiency due to the need to learn singulation and grasping simultaneously. Alternatively, a two-stage method trains separate policies for singulation and grasping, but it leads to longer execution times and suboptimal transitions, as it fails to leverage their synergy.
@@
Detailed definitions of these rewards can be found in Tab. I, where the subscript $t$ is omitted for simplicity. The rewards are defined as follows: $r_{t}^{P}$ encourages the palm to approach the target object; $r_{t}^{J}$ and $r_{t}^{F}$ promote grasp formation through joint and fingertip contacts; $r_{t}^{L}$ encourages lifting once contact is established; $r_{t}^{S}$ drives separation of the target from obstacles; $r_{t}^{G}$ guides motion toward the goal position; and $r_{t}^{B}$ provides a bonus for successful singulation and grasping. $d_{t}^{P}$ and $d_{t}^{J}$ denote the minimum distances from the palm and hand links to the target, respectively. $\{p_{i}^{\text{target}}\}_{i=1}^{N_{p}}$ are the target surface points; $p^{\text{palm}}$, $\{p_{j}^{\text{link}}\}_{j=1}^{m}$, and $\{p_{j}^{\text{fingertip}}\}_{j=1}^{h}$ denote the palm, link, and fingertip positions; $p^{\text{goal}}$ and $p^{\text{target}}$ are the goal and current positions of the target object; $\{p_{i}\}_{i=1}^{n}$ are the obstacle positions; and $a^{P_{t_{z}}}$ represents palm translation along +z+z, corresponding to the lifting motion.
The reward function consists of two main components: an approach reward that encourages the hand to move toward the target object and a lifting reward that promotes object elevation after contact is established. To further facilitate singulation, the singulation reward $r_{t}^{S}$ is incorporated into both components, incentivizing the hand to separate the target object from surrounding obstacles. The transition between these two reward stages is achieved by a contact criterion specified by $d_{t}^{P}<0.06$ and $d_{t}^{J}<0.2$. These values are selected such that the palm and fingers are close enough for grasping. The reward transitions from approaching to grasping, ensuring a smooth progression toward a successful grasp.
Despite this unified learning framework, training remains highly challenging due to the increasing complexity of clutters. To further improve learning efficiency and policy generalization, we introduce Clutter Arrangement Curriculum Learning, which progressively increases clutter complexity.
@@ III-B Clutter Arrangement Curriculum Learning
To ensure successful and efficient object singulation and grasping in SNCs, we begin by training the teacher policy with privileged information from the simulation and adopt clutter arrangement curriculum learning, allowing our teacher policy to progressively improve as object diversity and spatial complexity increase. Using Proximal Policy Optimization (PPO) [40], we optimize the policy to maximize the cumulative discounted reward $E[\sum_{t=1}^{T}\gamma^{t-1}r_{t}]$, enabling effective RL.
III-B1 Clutter Generation
To facilitate the training and testing of policies across various object configurations, we introduce a clutter generation module designed to create diverse object-based tasks. Our clutters consist of cuboid-shaped objects with varying quantities (from 4 to 8) and shapes (1×1, 1×2, and 1×3 cuboids). The generation for two types of clutters are introduced below, as shown in Fig. 3.
Dense Arrangements Generation. This type of task arranges different quantities of the surrounding 1×1 cuboids densely near the target object to create an extreme scenario that challenges the singulation and grasping policies under dense and narrow conditions.
Random Arrangements Generation. This type of task arranges objects of different quantities and shapes randomly around the target object for grasping, mainly to test the generalization of the singulation and grasping policies.
For simplicity, we use D/R-$n$ to denote task setting with $n$ objects for dense (D) or random (R) arrangements, such as D-8. In our curriculum, only D-8 and R-8 are SNC tasks, while D-4/6 and R-4/6 are intermediate stages to aid learning and are not necessarily SNCs.